An object is placed 14.5 cm in front of a convex m

What is the magnification of the object in the previous problem?
Answer #1

1/f = 1/u+1/v

-1/24.5 = 1/14.5 + 1/v

1/v = -1/24.5-1/14.5

v = 9.1 cm

MAGNIFICATION = v/u = 9.1/14.5 = 0.628